OpenStudy (andriod09):

how do you do Pythagorean Theory, I used to know how, I forgot becuase I haven't used it in forever. I have \[a^2+b^2=c^2\] \[a^2+15^2=12^2\] \[a^2+255=144\] i don't know where to go after this. :/

OpenStudy (anonymous):

What exactly is it that you're trying to do?

OpenStudy (andriod09):

i'm trying to solve this: 

OpenStudy (anonymous):

Okay. Pythagoras says that "the square of the hypotenuse is equal to the sum of the squares of the sides". The hypotenuse of a triangle is the longest side. The "15" in your diagram is the hypotenuse. In your equation a^2+b^2=c^2, "c" represents your hypotenuse, and you are looking for the length of "a" Here's what you do: a = sqrt(15^2-12^2)

OpenStudy (andriod09):

yes, i know this. i don't know where to go after what i have. i have it all multiplied out, but i'm stuck there, i have this so far: \[a^2+255=144\] which is \[a^2+15^2=12^2\] multiplied out.

OpenStudy (anonymous):

No. a^2 = 15^2-12^2 a^2 = 225-144

OpenStudy (anonymous):

a small hint: when two squares add, you'd always get a bigger number.. never smaller

OpenStudy (andriod09):

again, i haven't done this in years dude. soz i'm a fail. i should have known this.

OpenStudy (anonymous):

S'okay...we're here to remind you... :-)

OpenStudy (anonymous):

@qweqwe123123123123111 lets keep imaginaries outta here :P

OpenStudy (anonymous):